OpenStudy (anonymous):

HELP PLEASE!! MEDALS WILL BE GIVEN!! Express the solutions for x as Logarithms: 10^2x-5(10^x) +6 = 0

OpenStudy (anonymous):

Maybe a substitution will clear things up: let \(y=10^x\): \[10^{2x}-5(10^x)+6=0~~\iff~~y^2-5y+6=0\]

OpenStudy (anonymous):

yea but i have to change it into logarithmic form

OpenStudy (anonymous):

actually nm...i got it thank you :)

OpenStudy (anonymous):

yw